# Towards Zero-Shot Interpretable Human Recognition: A 2D-3D Registration Framework

### v1

@@ 1 Introduction
It is known that the remarkable ability of humans to recognize objects relies in prototypes somehow stored in our brain and matched to a compressed version of the input data, pre-processed by the visual system. For example, when looking at a photograph of an individual ’A’ whom we know, we can compare it with his 3D representation (prototype), justifying whether both correspond or not: ”those are the arms and legs thcikness and length of individual ’A’, and the head shape is also the same”. This process is not only the key for recognition, but also accounts for explainability by pointing out which parts of the image correspond to the prototype, as in a registration process.
@@
As a response to the above outlined difficulties, one of the our constraints from the beginning was to rely exclusively in synthetic data for learning purposes. This allows to generate a practically unlimited number of images, capturing necessary variations without the logistical limitations associated to collecting extensive real data. Surely, even though we train the model using synthetic dataset, all the test experiments were carried in real data, highlighting its zero-shot learning capability [7]. In our context, we use this term to express the model’s ability to generalize to data modalities it hasn’t seen during training. The rationale behind this lies in the capability to render high-quality data from just a 3D mesh of an individual. This enables the model to recognize individuals in entirely different conditions, such as real-world scenarios.
@@ 2.2 Interpretable Object Recognition
The post hoc methods refers to the ability to interpret or explain the decisions made by a machine learning model after it has been trained [22]. In this group, methods such as LIME and SHAP stand out [17, 12]. The former works by perturbing instances around a data point, observing original model predictions, and fitting a simpler model (e.g., linear regression) to explain the original model’s behavior locally. The latter explains the model predictions by attributing the impact of each input feature, using principles from game theory.
@@ 3 Proposed Method
Our proposed method is divided into two main phases: generative phase and learning phase. Fig. 2 illustrates the workflow of all the proposed method phases.
@@
Figure 2: Cohesive perspective of the whole pipeline proposed in this paper: In the initial phase, we compile highly detailed 3D meshes of each individual and use them to generate images with a wide range of factors. In the subsequent phase, the model learns to semantically match the individual images with their 3D representation (point cloud).
@@ 3.3 Generative Phase
Firstly, we extract the SMPL [10] mesh of the individual using SMPLify [1]. This type of mesh allows obtaining a 3D representation of the individual’s body that is very close to their real physique. However, the SMPL [10] mesh does not preserve details of the individual’s head, hair, and face. For this reason, we extract a detailed mesh of the head and replace the head of the SMPL [10] mesh with this one. In this way, we obtain a highly detailed and complete 3D representation of the individual.
@@ 3.4 Framework Architecture
The encoders in our model use a basic VGG [19] style, with each block used in our network consisting primarily of a sequence of convolutional layers with a (3x1x1) configuration, followed by batch normalization and ReLU activation like in [4].
@@ 4.2 Qualitative Results
While the model was trained using synthetic data, it is more meaningful to evaluate the its performance with real-world data. In this way, we tested the model with all the real images provided by RenderPeople [16]. For each of these images, the respective 3D representation was passed through. Figure 4 shows some examples of the obtained results, providing a general overview of the model’s performance. Here, the colored regions symbolize the 2D-3D semantic registration, representing the region of image pixels that have semantic correspondences with the vertices of the point cloud.
@@ 4.3 Quantitative Results
Fig. 5 displays the confusion matrices of correspondence rates varying the threshold parameter, denoted as $t$. This parameter allows for adjusting the confidence of the obtained correspondences. We observe that a higher value of $t$ results in fewer false positives, emphasizing the diagonal elements of the matrices. Consequently, the maximum value of $\rho$ decreases as the quantity of correspondences also decreases.

### v2

@@ 1 Introduction
It is known that the remarkable ability of humans to recognize objects relies on prototypes somehow stored in our brain and matched to a compressed version of the input data, pre-processed by the visual system. For example, when looking at a photograph of an individual ’A’ whom we know, we can compare it with his 3D representation (prototype), justifying whether both correspond or not: ”those are the arms and legs thickness and length of individual ’A’, and the head shape is also the same”. This process is not only the key for recognition, but also accounts for explainability by pointing out which parts of the image correspond to the prototype, as in a registration process.
@@
As a response to the above outlined difficulties, one of our constraints from the beginning was to rely exclusively in synthetic data for learning purposes. This allows to generate a practically unlimited number of images, capturing necessary variations without the logistical limitations associated to collecting extensive real data. Surely, even though we train the model using synthetic dataset, all the test experiments were carried in real data, highlighting its zero-shot learning capability [10]. In our context, we use this term to express the model’s ability to generalize to data modalities it hasn’t seen during training. The rationale behind this lies in the capability to render high-quality data from just a 3D mesh of an individual. This enables the model to recognize individuals in entirely different conditions, such as real-world scenarios.
@@ 2 Related Work
2.2 Semantic Correspondence
Recent advancements in semantic correspondence have significantly improved the matching of similar objects and objects parts across images using deep learning techniques.
SFNet [11] uses binary foreground masks to guide a CNN in learning dense flow fields between image instances. The method emphasizes object-aware learning to reduce background clutter and employs a novel kernel soft argmax layer for accurate matching.
Deep ViT [2] utilizes deep features from Vision Transformers (ViTs), specifically from a self-supervised DINO-ViT [4] model. These features encode high-level semantic information and enable robust zero-shot applications in co-segmentation and semantic correspondences without extensive training, showing competitive performance against supervised methods.
Neural Best-Buddies [1] introduces a method for identifying sparse correspondences between images with semantically related parts. Utilizing a coarse-to-fine approach with deep features from a pre-trained CNN, this method finds mutual nearest neighbors at various abstraction levels and refines them to handle cross-domain variations.
@@ 2.3 Interpretable Object Recognition
The post hoc methods refer to the ability to interpret or explain the decisions made by a machine learning model after it has been trained [26]. In this group, methods such as LIME and SHAP stand out [21, 16]. The former works by perturbing instances around a data point, observing original model predictions, and fitting a simpler model (e.g., linear regression) to explain the original model’s behavior locally. The latter explains the model predictions by attributing the impact of each input feature, using principles from game theory.
@@ 3 Proposed Method
Our proposed method is divided into three phases: generative phase, learning phase and inference phase. Fig. 2 illustrates the workflow of the generative and learning phases.
@@
Figure 2: Cohesive perspective of the whole pipeline proposed in this paper: In the initial phase, we compile highly detailed 3D meshes of each individual and use them to generate images with a wide range of factors. In the subsequent phase, the model learns to semantically match the individual images with their 3D representation (point cloud) by pulling similar features together and pushing dissimilar features apart.
@@ 3.3 Generative Phase
Firstly, we extract the SMPL [14] mesh of the individual using SMPLify [3]. This type of mesh allows obtaining a 3D representation of the individual’s body that is very close to their real physique. However, the SMPL [14] mesh does not preserve details of the individual’s head, hair, and face. For this reason, we extract a detailed mesh of the head and replace the head of the SMPL [14] mesh with this one. In this way, we obtain a highly detailed and complete 3D representation of the individual. Further details on the extraction of the detailed mesh of head from each individual, and the process of replacing them on the 3D mesh generated through SMPL [14], are described in the Subsection 4.1.
@@ 3.4 Framework Architecture
The encoders in our model use a basic VGG [23] style, with each block used in our network consisting primarily of a sequence of convolutional layers with a ($3\times 1\times 1$) configuration, followed by batch normalization and ReLU activation like in [7].
@@ 4 Experiments and Discussion
In this section, we present the results obtained from our model after it was trained on the synthetic dataset and tested on the real images test set. We discuss the dataset rationale for 3D human generation and some implementation details. To perform a detailed analysis of the model’s performance, we present and discuss both qualitative and quantitative results. Finally, we highlight some notable cases that are particularly worth mentioning.
@@ 4.2 Qualitative Results
While the model was trained using synthetic data, it is more meaningful to evaluate its performance with real-world data. In this way, we tested the model with all the real images provided by RenderPeople [20]. For each of these images, the respective 3D representation was passed through. Figure 4 shows some examples of the obtained results, providing a general overview of the model’s performance. Here, the colored regions symbolize the 2D-3D semantic registration, representing the region of image pixels that have semantic correspondences with the vertices of the point cloud.
@@ 4.3 Quantitative Results
Fig. 5 displays the confusion matrices of correspondence rates varying the threshold parameter, denoted as $t$. This parameter allows for adjusting the confidence of the obtained correspondences. We observe that a higher value of $t$ results in fewer false positives, emphasizing the diagonal elements of the matrices. Consequently, the maximum value of $\rho$ decreases as the quantity of correspondences also decreases. We can have a more objective idea by observing the F1-Scores of each class for $t=0.9$, by examining the Table 1.
Class Name F1-Score (0.9)
005 0.839
010 0.434
047 0.231
069 0.502
070 0.646
009 0.478
026 0.440
055 0.220
061 0.217
091 0.406
Table 1: F1-Scores per Class with $t=0.9$.